OpenStudy (thatcluelessgirl):

need help medal + fan Write the equation of a line in slope intercept form that is perpendicular to the line y = –4x and passes through the point (2, 6).

OpenStudy (tkhunny):

Please write out the Slope-Intercept form Please show YOUR work.

OpenStudy (thatcluelessgirl):

I have no clue do I put a dot on negative 4 y and go up 2 over six

OpenStudy (tkhunny):

Why are you describing a graphing procedure? Simply write down the Slope-Intercept Form. This is your first step toward a solution.

rishavraj (rishavraj):

see uu got the slope of one line.... now if two lines are perpendicular..then product of their slopes is -1... using tht u can find the slope of other line and proceed

OpenStudy (thatcluelessgirl):

@tkhunny were do I get the second line

OpenStudy (tkhunny):

You didn't do what I asked. You get the second line by WRITING OUT the Point-Slope form and then supplying it with the necessary parameters. In this case, you are given the slope indirectly. You will be required to know, on an exam, how perpendicular lines relate to each other as far as their slopes are concerned.

OpenStudy (triciaal):

different approach given the line Y = -4x knowing y = mx + b can you identify the slope m? what do you know about perpendicular lines? The product of the slopes is -1 rephrased if the slope of a line is m the slope of the perpendicular is -1/m y = mx + b given (x,y) = (2,6) and m you calculated can you find b now rewrite for any (x,y)

OpenStudy (thatcluelessgirl):

I'm so confused

OpenStudy (thatcluelessgirl):

how would I set it up

OpenStudy (triciaal):

read my response slower and let me know where you are getting stuck so I can clear it up

OpenStudy (tkhunny):

WRITE OUT the Point Slope form.

OpenStudy (thatcluelessgirl):

knowing y = mx + b can you identify the slope m?

OpenStudy (thatcluelessgirl):

what do I put in for mx +b

OpenStudy (triciaal):